Show the details clearly please.
The problem from chapter 4
media%2F538%2F538a48d4-f5f2-4327-a450-85
Answer #1

Ans:

Exponential distribution has memoryless property.

So,

P(T<=15/T>10)=P(T<=5)

P(T<=5)=1-e-0.2*5

=1-e-1=0.6321